# Cost-benefit analysis of an AI-driven operational digital platform for integrated electric mobility, renewable energy, and grid management

### v1

@@ 5 Discussion
The comprehensive CBA conducted for the AI-driven ODP clearly indicates that the proposed platform offers substantial net benefits for the integrated management of electric mobility, renewable energy, and grid operations. The aggregated positive net present value (NPV) of €356.7 million and a benefit-cost ratio (BCR) of 1.41 provide a strong economic rationale, signifying that every Euro invested generates approximately €1.41 in benefits over the 10-year analysis period.
@@
The case studies across Austria, Hungary, and Slovenia demonstrate the ODP’s adaptability and value proposition across varied national energy and transport landscapes. The substantial reliability benefits (€348 million in GSMS, alongside €171 million in CSDR-PLR as an economic benefit) underscore the ODP’s potential as a highly effective non-wires alternative, achieved through improved system visibility, AI-driven proactive control, and enhanced coordination across sectors. From a societal and policy perspective, the significant environmental benefits (€127 million from CO2 and air pollutant reductions) are critical, highlighting the ODP’s role in accelerating decarbonization. The robustness of AI models to real-world conditions such as data drift, concept drift, and unexpected events is planned to be managed through continuous performance monitoring, automated retraining pipelines, and robust data validation, which are all factored into the recurring OPEX (A.3), ensuring sustained long-term benefits.
@@
While the comprehensive CBA provides a strong economic rationale, several limitations warrant acknowledgment. Firstly, while rigorous backtesting and validation were performed, and simulations incorporate realistic assumptions, there remains a potential for variance in real-world AI model performance versus simulated benchmarks, particularly under unforeseen extreme conditions. Future pilot projects are essential for further empirical validation. Secondly, the optimal functioning of the ODP’s AI engine relies on sustained, high-quality, and comprehensive data availability from a multitude of disparate sources. These include real-time EV and ET fleet telemetry, smart charging infrastructure, renewable energy forecasts, market data from EPEX Spot and ENTSO-E, weather models, and public datasets from Eurostat and the EEA.
Thirdly, certain intangible benefits, such as enhanced user convenience, improved public health outcomes from cleaner air that are difficult to fully monetize, or the empowerment of consumers in energy management, were not explicitly quantified in monetary terms, potentially leading to an underestimation of the true societal value. Additionally, the study makes assumptions regarding rational user behavior in response to ODP incentives and information, which may vary in practice. The dynamic nature of the regulatory landscape influencing energy markets, data sharing, and cross-border energy flows introduces inherent uncertainties. Finally, while the three selected EU member states represent a diverse range, there are inherent challenges in quantitatively generalizing the precise findings to contexts vastly different from those studied, although the qualitative benefits of integrated, AI-driven energy and transport management are broadly applicable globally.
@@ 6 Conclusion
This paper has presented a comprehensive cost-benefit analysis of an AI-driven operational digital platform developed for the integrated management of electric vehicles, electric trucks, renewable energy sources, and the power grid. The CBA framework was designed in alignment with European Commission guidelines for energy infrastructure projects and tailored to explicitly capture AI-specific value streams. These include improvements in forecasting accuracy, operational optimization, grid reliability, and emissions reduction, each monetized using transparent and reproducible methodologies. Importantly, the framework incorporates both capital and operational expenditures, particularly those linked to AI model development, deployment, and maintenance, alongside benefits categorized as economic, reliability, and environmental.
@@
The analysis demonstrates that the proposed AI-driven ODP yields substantial net economic and societal value. With a projected net present value of approximately €356.7 million and a benefit-cost ratio of 1.41 over the 2026–2035 operational period, the investment case is strong. These results are driven by the ODP’s multi-layered architecture, where the AI core, featuring LSTM-based forecasting models, MILP-based optimization algorithms, and electric vehicle routing solvers, serves as a critical value enabler. These technologies jointly support enhanced energy and operational efficiency, predictive congestion mitigation, optimized demand response, and battery degradation reduction through health-aware charging strategies. The robustness of these benefits was confirmed through extensive stress-testing, including sensitivity analyses and Monte Carlo simulations.
@@
To validate cross-border applicability, the CBA was applied to Austria, Hungary, and Slovenia, each reflecting different energy and transport profiles. The findings illustrate how national characteristics influence the distribution of benefits: Austria achieves substantial fleet efficiency gains and extended battery life; Hungary benefits from improved RES integration and reduced curtailment; and Slovenia sees value in grid flexibility and optimized ET logistics. These variations underscore the scalability and adaptability of the ODP to diverse regulatory, technical, and market contexts.
@@
Beyond financial metrics, the study offers strategic insights. For policymakers, the results affirm the role of advanced digital infrastructure in achieving EU-wide decarbonization goals under REPowerEU and Fit for 55. The ODP presents a viable pathway to accelerate transport electrification, enhance grid reliability, and reduce emissions through AI-based coordination. For industry stakeholders, including DSOs, TSOs, fleet operators, and aggregators, the analysis provides a quantitative basis for investing in interoperable, intelligent systems capable of delivering both operational gains and new service opportunities. At the same time, the study acknowledges key limitations. The optimal functioning of the AI engine depends on sustained access to high-quality, heterogeneous data sources, such as real-time telemetry from EVs and ETs, smart charging infrastructure, market platforms (e.g., EPEX Spot, ENTSO-E), and weather models. Moreover, generalizing the results to other countries requires careful consideration of differences in data infrastructure, policy environments, and electrification maturity. Additionally, intangible benefits, such as improved public health from air quality gains or increased user convenience, were not monetized, likely leading to a conservative estimate of the platform’s true societal value.
@@
In conclusion, the integrated management of EVs, ETs, RES, and the grid via AI-driven digital platforms is shown to be not only technically feasible but also economically and environmentally beneficial. The proposed ODP offers a compelling blueprint for next-generation energy–mobility convergence, supporting the emergence of smart, flexible, and low-carbon energy ecosystems. By holistically orchestrating distributed assets through AI, the platform bridges the gap between high-level policy goals and on-the-ground operational needs.
@@
Future work should focus on several key areas to further validate and expand the ODP’s capabilities and impact. This includes the initiation of pilot projects for empirical validation of the ODP’s AI models and overall system performance in real-world operational environments, providing concrete data on benefits and operational challenges. A deeper exploration of advanced V2X (Vehicle-to-Everything) services, including vehicle-to-home (V2H) and vehicle-to-grid (V2G) applications beyond ancillary services, could unlock further synergies. Detailed distribution grid impact modeling, specifically for local voltage stability and transformer loading under high EV penetration and dynamic charging, is crucial. Research into advanced cybersecurity measures tailored for AI-driven energy platforms, including AI-specific threat detection and mitigation, is paramount given the critical infrastructure nature. Policy and market design innovation is essential to incentivize broad stakeholder participation and ensure fair value distribution across the ecosystem. Finally, extending the ODP concept to integrate other flexible resources like smart heat pumps, stationary battery storage, and industrial flexible loads could further enhance system-wide optimization and accelerate decarbonization efforts.
@@ Appendix A Detailed cost-benefit analysis methodology and data
This appendix provides further details on the methodologies and data underpinning the CBA presented in the main paper, ensuring transparency and reproducibility of the quantitative analysis.
@@ A.1 Benefit monetization formulas and key parameters
This section details the primary formulas and key parameters used for monetizing benefits in Step 4: Monetize benefits of the CBA, as discussed in Section 4. All monetary values are expressed in real terms (Euro 2025), and future cash flows are discounted to their present value using a social discount rate of . Each benefit’s realization is linked to the capabilities of the AI engine described in Section 3.1. In the following subsections, we present the formulas for calculating each benefit, along with the specific country-related assumptions, technological assumptions, and projected data.
@@ A.1.1 Reduction in operational downtime (ROD) cost
where is the baseline operational downtime cost222A conservative 24.2 million €/year for the three countries is taken in our calculations., is the energy self-efficiency improvement fraction333For this benefit calculation we assume and , and is the ODP-enabled downtime cost savings, reflecting faster fault recovery and improved grid reliability due to AI-driven predictive maintenance and situational awareness (Section 3.1). The product captures the adjusted downtime cost under improved self-efficiency, and multiplying by converts this into the annual benefit from avoided unserved energy, consistent with analogous smart-grid reliability improvements [53].
@@ A.1.2 Revenue from optimized energy trading and ancillary services (ROETAS)
This benefit quantifies increased revenues and cost savings from optimized market participation due to the ODP’s AI forecasting and optimization capabilities.
where stands for the energy arbitrage revenue at hour under an ODP-enabled system and is the corresponding revenue without ODP. The hourly arbitrage revenue itself is calculated as where is the ODP-enabled flexible capacity (MWh), is the AI-identified price spread (€/MWh), and is the round-trip efficiency. To explicitly consider the efficiency improvement on RES, ET, and EV due to ODP444For our calculation, the EV/ET efficiency improvement due to ODP is assumed to be 5%, of total annual energy, and €/MWh. Moreover, the RES capacity factor CF represents typical full-load hours for renewables; we use a conservative value of 1,200 h/year., one may replace with where denotes the net uplift fraction and [€/MWh] is the assumed average electricity price. The term represents the annual baseline renewable-energy output, defined as , where [MW] is the installed RES capacity and [h/yr] is its full-load hours. Similarly, is the annual baseline EV + ET energy in MWh, with [MW] the aggregated EV + ET charging capacity and h/yr. The set lists all ancillary-service products. For each and each hour , [€/MWh] and [€/MWh] denote the ODP’s revenue and cost for service , respectively, while [€/MWh] and [€/WWh] denote the corresponding baseline values under non-ODP operation.
A.1.3 Cost savings from demand response and peak load reduction (CSDR-PLR)
Avoided costs from reducing system peak demand are realized through ODP’s AI-driven demand response (DR) and peak-load-shifting capabilities. We model the annual benefit as
Where the first term is benefits from cost savings from demand response, and the second term is benefits from peakload revenue. [€/MWh] is the unit cost saving from demand response (CSDR) per shifted MWh; [MW] is the installed RES capacity and [h/yr] its full-load hours, hence is the annual RES-generation baseline in MWh. The factor is the fraction of RES output that can be shifted via DR555, €/MWh, and in our calculations.. Similarly, [MW] is the aggregated EV charging capacity, [MW] is the total ET charging capacity, the terms and are the annual EV and ET baseline energy (MWh) multiplied by the shiftable fractions and .
In the second term, [€/MWh] denotes the peak-load reduction (PLR) revenue per shifted MWh, and is the fraction of total annual energy that can be shifted away from peak hours. We also have:
where is the annual volume (MWh) shifted out of peak hours. Thus, the second term represents revenue from peak-load reduction at peak-period prices.
@@ A.1.4 Avoided energy curtailment (AEC)
Monetized value of RES utilized rather than curtailed, enabled by ODP’s AI-driven RES forecasting and flexible load matching. This benefit is calculated using the following formula:
where is the fraction of RES curtailment avoided666=1% and 50 €/MWh in our calculation., [MW] is the installed RES capacity, and [€/MWh] is the unit cost associated with avoided curtailment. Thus, gives the annual baseline RES generation (MWh), multiplying by yields the avoided MWh, and multiplying by converts to euros.
@@ A.1.5 Fleet energy and operational savings (FES)
Reduced EV/ET operational costs are achieved through ODP’s AI-optimized routing and charging, which extend battery lifespan and improve energy efficiency. The total annual benefit is calculated as follows:
where is the overall ODP-enabled savings fraction777We consider 1%, 800 €/year, 1920 €/year in our calculation., and are the total number of EVs and ETs, respectively, and are the effective efficiency factors capturing reduced energy consumption and extended battery life, and and are the annual per-vehicle energy and battery-degradation savings under ODP. The factor consolidates multiple AI-driven improvements: battery-health optimization (reducing degradation by via SoC and C-rate management [36, 37, 38]) and routing efficiency gains (3–5% savings from E-VRPTW optimization). Thus, reflects a conservative 1% net reduction in total fleet operating cost when multiplying the baseline EV and ET savings terms.

### v2

@@ 5 Discussion
5.1 Key findings and insights
The proposed ODP–CBA framework is illustrated with case studies from three countries, but it is general and scalable beyond these settings. While the underlying methodology remains unchanged, differences in sensor unit costs or institutional settings would shift the numerical results but not the validity of the framework itself. The framework’s modular architecture centered on EVs, ETs, RES, and the grid, which remains valid for any jurisdiction, provided basic assumptions and data are available. Scaling to larger systems (e.g., Germany) primarily requires adjusting inputs (fleet sizes, RES capacities, growth projections, grid complexity, sensor unit costs), which change the magnitude of costs and benefits but not the methodology. This design supports both computational scalability and institutional transferability. To this end, the comprehensive CBA conducted for the AI-driven ODP clearly indicates that the proposed platform offers substantial net benefits for the integrated management of electric mobility, renewable energy, and grid operations. The aggregated positive NPV of €356.7 million and a BCR of 1.41 provide a strong economic rationale, signifying that every Euro invested generates approximately €1.41 in benefits over the 10-year analysis period.
@@ 5.1 Key findings and insights
The case studies across Austria, Hungary, and Slovenia demonstrate the ODP’s adaptability and value proposition across varied national energy and transport landscapes. The substantial reliability benefits (€348 million in GSMS, alongside €171 million in CSDR-PLR as an economic benefit) underscore the ODP’s potential as a highly effective non-wires alternative, achieved through improved system visibility, AI-driven proactive control, and enhanced coordination across sectors. From a societal and policy perspective, the significant environmental benefits (€127 million from CO2 and air pollutant reductions) are critical, highlighting the ODP’s role in accelerating decarbonization. The robustness of AI models to real-world conditions, such as data drift, concept drift, and unexpected events, is planned to be managed through continuous performance monitoring, automated retraining pipelines, and robust data validation, all of which are factored into the recurring OPEX (A.3), ensuring sustained long-term benefits.
In addition to the aggregated results, Table 9 presents the country-specific CBA outcomes and Monte Carlo simulations for Austria, Hungary, and Slovenia. This disaggregation highlights national variations in benefits and costs due to grid structure, RES penetration, and policy context. The results confirm that economic benefits are the dominant category in all three countries, while reliability and environmental benefits contribute proportionally more in Hungary and Slovenia. Across all countries, the benefit–cost ratio remains above one under the 90% confidence interval, demonstrating the robustness of the investment case even under country-specific risk and uncertainty.
5.2 Limitations and implementation challenges
@@
While the comprehensive CBA provides a strong economic rationale, several limitations warrant acknowledgment. First, while rigorous validation were performed, and simulations incorporate realistic assumptions, there remains a potential for variance in real-world AI model performance versus simulated benchmarks, particularly under unforeseen extreme conditions. Future pilot projects are essential for further empirical validation. Second, the optimal functioning of the ODP’s AI engine relies on sustained, high-quality, and comprehensive data availability from a multitude of disparate sources. These include real-time EV and ET fleet telemetry, smart charging infrastructure, renewable energy forecasts, market data from EPEX Spot and ENTSO-E, weather models, and public datasets from Eurostat and the EEA. Third, our CBA does not include primary behavioral studies (e.g., discrete-choice experiments or user trials), beyond stakeholder engagement through interviews and workshops. Although behavioral factors are incorporated indirectly via the stakeholder-informed multi-criteria assessment and sensitivity analyses on adoption and usage, targeted behavioral studies to estimate adoption elasticities, routing/charging preferences, and data-sharing incentives would further strengthen the evidentiary basis and validity of the CBA. Four, Although a 10-year window is appropriate for a digital platform with sub-decadal upgrade cycles, some grid-adjacent hardware may realize benefits over longer periods. Evaluating selected components over 20–25 years—and reporting horizon sensitivities—can further validate long-run value in infrastructure-heavy contexts.
In addition, the following three prevalent precautions should be considered further considered:
Valuation and behavioral uncertainty: Some societal benefits (e.g., improved air quality, user convenience, consumer empowerment) are not fully monetized, which may understate ODP’s total value. In addition, the analysis assumes rational uptake of incentives by EV/ET users and fleets; however, real-world participation and rebound effects may differ, necessitating the use of behavioral modeling and longitudinal evidence.
Technical robustness and data risk: Large-scale data exchange and cloud operation introduce cybersecurity and governance risks. Model accuracy can also drift as mobility patterns, RES penetration, and technology costs evolve. Mitigation requires strong data governance frameworks (including privacy-preserving analytics and independent audits), continuous retraining and performance monitoring, and operational fallbacks. All these evaluations could be better estimated if there were already operating ODPs, even on a small scale.
Regulatory and institutional fit: Scalability depends on alignment with grid codes, tariff design, and data-sharing rules. Fragmented or evolving policies can increase coordination costs and slow deployment; staged pilots and regulatory sandboxes can help mitigate implementation risks and harmonize standards. That said, qualitative stakeholder analyses may not fully capture behavioral and adoption challenges. Empirically estimating the costs and benefits of these elements during implementation could enhance the reliability of the CBA framework.
@@ 6 Conclusion
This paper presents a comprehensive cost-benefit analysis of an AI-driven operational digital platform developed for the integrated management of electric vehicles, electric trucks, renewable energy sources, and the power grid. The CBA framework was designed in alignment with European Commission guidelines for energy infrastructure projects and tailored to explicitly capture AI-specific value streams. These include improvements in forecasting accuracy, operational optimization, grid reliability, and emissions reduction, each of which is monetized using transparent and reproducible methodologies. Importantly, the framework incorporates both capital and operational expenditures, particularly those linked to AI model development, deployment, and maintenance, alongside benefits categorized as economic, reliability, and environmental.
@@
The analysis shows that the proposed AI-driven ODP delivers substantial economic and financial value. Over the 2026–2035 operating period, it achieves a projected net present value of about €356.7 million and a benefit–cost ratio of 1.41, indicating a strong investment case. These outcomes are driven by the platform’s multi-layered architecture; its AI core—combining optimization and machine-learning algorithms—is a critical value enabler. The ODP improves energy and operational efficiency, enables predictive congestion mitigation, optimizes demand response, and reduces battery degradation through health-aware charging strategies. Using a dedicated framework, the CBA systematically monetized both benefits and investment/operating costs. Robustness was confirmed through sensitivity analyses and Monte Carlo simulations.
@@
To validate cross-border and cross-sector applicability, the CBA was applied to Austria, Hungary, and Slovenia—each with distinct energy and transport profiles. The results illustrate how national characteristics shape the distribution of benefits. At the country level, Austria achieves substantial fleet-efficiency gains and extended battery-life benefits; Hungary sees improved RES integration and reduced curtailment; and Slovenia has a comparative advantage in grid flexibility and optimized ET logistics. These variations underscore the ODP’s scalability and adaptability across diverse regulatory, technical, and market contexts.
@@
Beyond financial metrics, the study offers strategic insights. For policymakers, the results confirm the crucial role of advanced digital infrastructure in achieving EU-wide decarbonization goals under REPowerEU and the Fit for 55 initiatives. The ODP presents a viable pathway to accelerate transport electrification, enhance grid reliability, and reduce emissions through AI-based coordination. For industry stakeholders, including DSOs, TSOs, fleet operators, and aggregators, the analysis provides a quantitative basis for investing in interoperable, intelligent systems capable of delivering both operational gains and new service opportunities.
At the same time, the study acknowledges key limitations. The optimal functioning of the AI engine depends on sustained access to high-quality, heterogeneous data sources, such as real-time telemetry from EVs and ETs, smart charging infrastructure, market platforms (e.g., EPEX Spot, ENTSO-E), and weather models. Moreover, generalizing the results to other countries requires careful consideration of differences in data infrastructure, policy environments, and electrification maturity. Additionally, intangible benefits, such as improved public health from air quality improvements or increased user convenience, were not monetized, likely resulting in a conservative estimate of the platform’s true societal value.
@@
In conclusion, the integrated management of EVs, ETs, RES, and the grid via AI-driven digital platforms is shown to be not only technically feasible but also economically and environmentally beneficial. The proposed ODP offers a compelling blueprint for next-generation energy–mobility convergence, supporting the emergence of smart, flexible, and low-carbon energy ecosystems. By holistically orchestrating distributed assets through AI models, the platform bridges the gap between high-level policy goals and on-the-ground operational needs.
@@
Future work should focus on several key areas to further validate and expand the ODP’s capabilities and impact. This includes the initiation of pilot projects for empirical validation of the ODP’s AI models and overall system performance in real-world operational environments, providing concrete data on benefits and operational challenges. A deeper exploration of advanced V2X (Vehicle-to-Everything) services, including vehicle-to-home (V2H) and vehicle-to-grid (V2G) applications, could unlock further synergies beyond ancillary services. Detailed distribution grid impact modeling, specifically for local voltage stability and transformer loading under high EV penetration and dynamic charging, is crucial. Research into advanced cybersecurity measures tailored for AI-driven energy platforms, including AI-specific threat detection and mitigation, is paramount given the critical infrastructure nature. Policy and market design innovation are essential to incentivize broad stakeholder participation and ensure fair value distribution across the ecosystem. Finally, extending the ODP concept to integrate other flexible resources such as smart heat pumps, stationary battery storage, and industrial flexible loads could further enhance system-wide optimization and accelerate decarbonization efforts.
@@ Appendix A Detailed cost-benefit analysis methodology and data
This appendix provides further details on the methodologies and data underpinning the CBA presented in this paper, ensuring transparency and reproducibility of the quantitative analysis.
@@ A.1 Benefit monetization formulas and key parameters
This section presents the primary formulas and key parameters used for monetizing benefits using CBA framework. All monetary values are expressed in real terms (Euro 2025), and future cash flows are discounted to their present value using an economic discount rate of 4%. Each benefit’s realization is linked to the capabilities of the AI engine described in the ODP architecture. In the following, we present the formulas for calculating each benefit.
@@ A.1.1 Reduction in operational downtime (ROD) cost
where is the baseline operational downtime cost, is the energy self‐efficiency improvement fraction, and is the ODP‐enabled downtime cost savings, reflecting faster fault recovery and improved grid reliability due to AI‐driven predictive maintenance and situational awareness. The product captures the adjusted downtime cost under improved self‐efficiency, and multiplying by converts this into the annual benefit from avoided unserved energy, consistent with analogous smart‐grid reliability improvements [U.S. Department of Energy, Office of Electricity, 2019]. Parameter values are presented in Table 10.
@@ A.1.2 Revenue from optimized energy trading and ancillary services (ROETAS)
This benefit captures the incremental revenues and cost savings enabled by the ODP’s AI forecasting and optimization across energy arbitrage and ancillary-service market:
where:
1.
and are the hourly energy–arbitrage revenues (in €) in hour with and without ODP, respectively. We compute hourly arbitrage revenue as
with the flexible energy shifted in hour (MWh), the realized buy/sell price spread (€/MWh), and the round-trip efficiency.
2.
denotes the set of ancillary-service products. For each and hour , and are the ODP’s hourly revenue and cost in €; and are the corresponding baseline values.
If one prefers a compact proxy for arbitrage gains due to efficiency/forecasting improvements, the first term in 2 can be approximated by
where is the net uplift fraction attributable to ODP (dimensionless), is the assumed average electricity price (€/MWh), is baseline annual renewable output (MWh) with installed capacity (MW) and full-load hours (h/yr), and is baseline annual EV+ET energy (MWh) given aggregate charging capacity (MW) and h/yr.
A.1.3 Cost savings from demand response and peak load reduction (CSDR‐PLR)
Avoided costs from reducing system peak demand are realized through ODP’s AI‐driven demand response (DR) and peak‐load‐shifting capabilities. We model the annual benefit as
where the first term represents the benefits from cost savingsresulting from demand response, and the second term represents the benefits from peakload revenue. [€/MWh] is the unit cost saving from demand response (CSDR) per shifted MWh; [MW] is the installed RES capacity and [h/yr] its full‐load hours, hence is the annual RES‐generation baseline in MWh. The factor is the fraction of RES output that can be shifted via ODP-enabled. Similarly, [MW] is the aggregated EV charging capacity, [MW] is the total ET charging capacity, the terms and are the annual EV and ET baseline energy (MWh) multiplied by the shiftable fractions and due to ODP.
In the second term, [€/MWh] denotes the peak‐load reduction (PLR) revenue per shifted MWh, and is the fraction of total annual energy that can be shifted away from peak hours. We also have:
where is the annual volume (MWh) shifted out of peak hours. Thus, the second term represents revenue from peak‐load reduction at peak‐period prices.
@@ A.1.4 Avoided energy curtailment (AEC)
Monetized value of RES utilized rather than curtailed, enabled by ODP’s AI‐driven RES forecasting and flexible load matching. This benefit is calculated using the following formula:
where is the fraction of RES curtailment avoided, [MW] is the installed RES capacity, and [€/MWh] is the unit cost associated with avoided curtailment. Thus, gives the annual baseline RES generation (MWh), multiplying by yields the avoided MWh, and multiplying by converts to euros.
@@ A.1.5 Fleet energy and operational savings (FES)
Reduced EV/ET operational costs are achieved through ODP’s AI‐optimized routing and charging, which extend battery lifespan and improve energy efficiency. The total annual benefit is calculated as follows:
where is the overall ODP‐enabled savings fraction, and are the total number of EVs and ETs, respectively, and are the effective efficiency factors capturing reduced energy consumption and extended battery life, and and are the annual per‐vehicle energy and battery‐degradation savings under ODP. The factor consolidates multiple AI‐driven improvements: battery‐health optimization (reducing degradation by 8–12% via SoC and C‐rate management [Maheshwari et al., 2020, Curry et al., 2023]) and routing efficiency gains (3–5% savings from E-VRPTW optimization). Thus, reflects a conservative 1% net reduction in total fleet operating cost when multiplying the baseline EV and ET savings terms.